OpenStudy (anonymous):

Help with EQUATIONS FOR FAN + MEDAL!!! Write an equation in point-slope form of the line that passes through (3, -5) and (2,3). Use the point (3,-5) in your equation.

OpenStudy (anonymous):

What is your question? Where the line passes through?

OpenStudy (daniellelovee):

you have to fit the points into y=mx+b

OpenStudy (daniellelovee):

sorry nvm I meant y-y1=m(x-x1) and then work PEMDAS backwards :)

OpenStudy (daniellelovee):

do you know how to do that? or need help?

OpenStudy (anonymous):

I might need some help @Daniellelovee

OpenStudy (daniellelovee):

ok I'll give you the equation and then you just find m(slope) from there

OpenStudy (daniellelovee):

y-y1=m(x-x1) 3-(-5)=m(2-3)

OpenStudy (daniellelovee):

and then just solve for m

OpenStudy (daniellelovee):

actually nvm they are just asking for the equation which I already wrote so

OpenStudy (daniellelovee):

but I would find the slope for the formula because some teachers are different and maybe your teacher wants you to find the slop and add it to y-y1=m(x-x1)

OpenStudy (anonymous):

okay thanks @Daniellelovee.

OpenStudy (daniellelovee):

your welcome